# Hallucinating 360°: Panoramic Street-View Generation via Local Scenes Diffusion and Probabilistic Prompting

### v2

@@ IV-A Experiments Details
Evaluation Metrics and Strategies. We employ multiple evaluation strategies to assess the effectiveness of our panorama generation method comprehensively. Specifically, we utilize Fréchet Inception Distance (FID) [47], Structural Similarity Index (SSIM) [46] for image quality assessment. Meanwhile, we also employ no-reference image quality assessment (i.e., BRISQUE [45], PIQE [49]) to evaluate the quality by comparing the generated images with ground-truth images and those from other methods. Additionally, we evaluated the images, assessing their IoU score for road segmentation and mean intersection over union on the pre-trained BEV semantic segmentation network. Furthermore, to more intuitively compare the performance of different methods, we ranked the results for each parameter and computed a ranking score.
Experiment Setting. As we are the first to propose the generation of panoramic images using stitched pinhole images, addressing the current challenges in image collection and annotation within the panoramic generation community, we modify the BEVControl [14] to serve as our baseline. Both the image and control signal are represented in 2D format (i.e., BEV Map, Mask Map, Depth Map). Meanwhile, we introduce more comparisons with the 3D generation model, Magicdrive [18]. Our experiments were conducted on two A6000 GPUs over a training period of ten days. We set the diffusion step to $20$ without using data augmentation strategies.
@@ V Ablation Studies
Settings BRISQUE [45$\downarrow$ PIQE [49$\downarrow$ FID [47$\downarrow$ Driv.$\uparrow$ Mean.$\uparrow$ Rank.$\downarrow$
@@
Table II: An in-depth analysis of LSDM. $\uparrow$/$\downarrow$ indicates whether a higher or lower value is preferred for each metric. “Mask” refers to the masking strategy, while “M.Crs.” denotes the cross-attention method built upon the masking strategy. Additionally, “Rota” represents the rotation method. “Driv.” refers to the segmentation results of drivable area, and “Mean” denotes the average performance across all BEV segmentation classes.
Settings BRISQUE [45$\downarrow$ PIQE [49$\downarrow$ FID [47$\downarrow$ Driv.$\uparrow$ Mean.$\uparrow$ Rank.$\downarrow$
@@ V-B Analysis of Different Prompts
To achieve more controllable panoramic image generation and avoid the labor-intensive task of manual panoramic image annotation, we leverage large language models to pre-annotate the stitched panoramic images, thereby generating imprecise yet informative control prompts. Furthermore, we propose the PPM, which enables robust panoramic generation. By incorporating depth information, PPM dynamically selects relevant image tokens and depth cues, surpassing the baseline in terms of image quality while significantly improving controllability. Additionally, we introduce a mask map, which, when combined with the image, depth map, and mask map, allows Percept360 to achieve state-of-the-art performance in overall scoring. With the joint contributions of PPM and LSMD, Percep360 achieves comprehensive improvements in both image quality and layout controllability.
@@ VI Conclusion and Future Work
Limitation and Future Work: In this paper, we focus on exploiting image coherence and the controllability of 2D features. The label mapping from six-view images to panoramic representations constitutes a critical component for efficient panoramic data generation. This will be addressed in our future work, where we intend to explore the transfer of pinhole-view annotations to panoramic domains. Furthermore, existing street-view generation approaches have largely overlooked the effective utilization of prompt features, which we identify as another key direction for future investigation.

### v3

@@ IV-A Experiments Details
Evaluation Metrics and Strategies. We employ multiple evaluation strategies to assess the effectiveness of our panorama generation method comprehensively. Specifically, we utilize Fréchet Inception Distance (FID) [10], Structural Similarity Index (SSIM) [34] for image quality assessment. Meanwhile, we also employ no-reference image quality assessment (i.e., BRISQUE [20], PIQE [21]) to evaluate the quality by comparing the generated images with ground-truth images and those from other methods. Additionally, we evaluated the images, assessing their IoU score for road segmentation and mean intersection over union on the pre-trained BEV semantic segmentation network. Furthermore, to more intuitively compare the performance of different methods, we ranked the results for each score and obtained a ranking score.
Experiment Setting. As we are the first to propose the generation of panoramic images using stitched pinhole images, addressing the current challenges in image collection and annotation within the panoramic generation community, we modify the BEVControl [44] to serve as our baseline. Both the image and control signal are represented in 2D format (i.e., BEV Map, Mask Map, Depth Map). Meanwhile, we introduce more comparisons with the 3D generation model, Magicdrive [6]. Our experiments were conducted on two A6000 GPUs over a training period of ten days. We set the diffusion step to $20$ without using data augmentation strategies. To accelerate the diffusion process, the image is processed at a resolution of (60, 1200) following [26].
@@ V Ablation Studies
[…]
@@
Table II: An in-depth analysis of LSDM. $\uparrow$$\downarrow$ indicates whether a higher or lower value is preferred for each metric. “Mask” refers to the masking strategy, while “M.Crs.” denotes the cross-attention method built upon the masking strategy. Additionally, “Rota” represents the rotation method. “Driv.” refers to the segmentation results of drivable area, and “Mean” denotes the average performance across all BEV segmentation classes.
[…]
@@ V-B Analysis of Different Prompts
To achieve more controllable panoramic image generation and avoid the labor-intensive task of manual panoramic image annotation, we leverage large language models to pre-annotate the stitched panoramic images, thereby generating imprecise yet informative control prompts. Furthermore, we propose the PPM, which enables robust panoramic generation. By incorporating depth information, PPM dynamically selects relevant image tokens and depth cues, surpassing the baseline in terms of image quality while significantly improving controllability. With the joint contributions of PPM and LSMD, Percep360 achieves comprehensive improvements in both image quality and layout controllability.
@@ VI Conclusion and Future Work
Limitation and Future Work: In this paper, we focus on exploiting image coherence and the controllability of 2D features. The label mapping from six-view images to panoramic representations constitutes a critical component for efficient panoramic data generation. Furthermore, existing street-view generation approaches have largely overlooked the effective utilization of prompt features, which we identify as another key direction for future investigation.